OpenStudy (anonymous):

A theater has 500 seats. Three-fourth of the seats are filled. How many seats are filled?

OpenStudy (anonymous):

375 seats are taken

OpenStudy (anonymous):

You there?

OpenStudy (anonymous):

why??

OpenStudy (anonymous):

because a little hint i learned is that when you wanna find %'s you multiply 500 by .75 to get 3-fourth's of 500

OpenStudy (gorv):

\[\frac{ 3 }{ 4 }*500=3*125=375\]

OpenStudy (anonymous):

or 500*.75=375

OpenStudy (anonymous):

\[500*.75=375 XD\]